OpenStudy (anonymous):

i have another one ready here goes, Manufactures of bird feeders cuts spacers from the tube thats 111/3 inches long. How many spacers can be cut from tube if spacer must be 2/3 inch thick

OpenStudy (anonymous):

\[(111/3)/(2/3) = (111/3)*(3/2)\]

OpenStudy (anonymous):

the threes cancel and you have 111/2, or 55 spacers and some extra

OpenStudy (anonymous):

no it was 11 1/3

OpenStudy (anonymous):

the fraction was separated

OpenStudy (anonymous):

oh lol so 11+(1/3) = 34/3 thus, 34/2 spacers can be made, 17